OpenStudy (anonymous):

Solve for x. |6x + 9| = 9 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

\[x=0\] \[x=-3\]

OpenStudy (anonymous):

Since it's an absolute value, that means x can be either a negative or positive. 6x+9=9 or -(6x+9)=9 Then you solve the equations from there. 6x+9=9 6x=0 x=0 -(6x+9)=9 -6x-9=9 -6x=18 x=-3